A spring of unstretched length l has a mass m with one end fixed to a rigid support .Assuming spring to be made of a uniform wire,find the kinetic energy possessed by it if its free end is pulled with uniform velocity v .
 
 

Grade:12

5 Answers

sarim khan
9 years ago
whenever u find such questions replace m of spring by m/3 ..now see how to do it.
k.e=1/2mv^2
the velocity of a point dx at a distance x from the rigid support will be vx/l.
let w=m/l(linear mass density)
so k.e of particle at x distance of dx thickness=1/2* w*dx*(vx/l)^2
int we get
k.e=1/2 *(m/3)*v^2
poopoo
7 years ago
fucking hate the way indians answer questions.... “replace m of spring by m/3” -like its sorcery
quote: “ 
whenever u find such questions replace m of spring by m/3 ..now see how to do it.
k.e=1/2mv^2
the velocity of a point dx at a distance x from the rigid support will be vx/l.
let w=m/l(linear mass density)
so k.e of particle at x distance of dx thickness=1/2* w*dx*(vx/l)^2
int we get
k.e=1/2 *(m/3)*v^2”
Moonshine
6 years ago

We can find the effective mass of the spring by finding its kinetic energy. This requires adding all the length elements' kinetic energy, and requires the following integral:

T =\int_m\frac{1}{2}u^2\,dm

Since the spring is uniform, dm=\left(\frac{dy}{L}\right)m, where L is the length of the spring. Hence,

=\frac{1}{2}\frac{m}{L}\int_0^L u^2\,dy T = \int_0^L\frac{1}{2}u^2\left(\frac{dy}{L}\right)m\!

The velocity of each mass element of the spring is directly proportional to its length, i.e. u=\frac{vy}{L}, from which it follows:

 

T =\frac{1}{2}\frac{m}{L}\int_0^L\left(\frac{vy}{L}\right)^2\,dy

 

=\frac{1}{2}\frac{m}{L^3}v^2\int_0^L y^2\,dy

 

=\frac{1}{2}\frac{m}{L^3}v^2\left[\frac{y^3}{3}\right]_0^L hence the answer =\frac{1}{2}\frac{m}{3}v^2
Samuel Garry
5 years ago

We can find the effective mass of the spring by finding its kinetic energy. This requires adding all the length elements' kinetic energy, and requires the following integral:

T =\int_m\frac{1}{2}u^2\,dm

Since the spring is uniform, dm=\left(\frac{dy}{L}\right)m, where L is the length of the spring. Hence,

=\frac{1}{2}\frac{m}{L}\int_0^L u^2\,dy T = \int_0^L\frac{1}{2}u^2\left(\frac{dy}{L}\right)m\!

The velocity of each mass element of the spring is directly proportional to its length, i.e. u=\frac{vy}{L}, from which it follows:

 

T =\frac{1}{2}\frac{m}{L}\int_0^L\left(\frac{vy}{L}\right)^2\,dy

 

=\frac{1}{2}\frac{m}{L^3}v^2\int_0^L y^2\,dy

 

=\frac{1}{2}\frac{m}{L^3}v^2\left[\frac{y^3}{3}\right]_0^L hence the answer =\frac{1}{2}\frac{m}{3}v^2
Samuel Garry
5 years ago
This is a very important concept of mass spring otherwise all the things done in our early classes of 11th is of mass Le spring but this is a very good question there is a very yes or no explanation of IT in mini box as this concept is widely used in olympiads level hope you understand about it